40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 342     Entry time: Wed Feb 27 22:05:03 2008
Author: John 
Type: Update 
Category: LSC 
Subject: Auxiliary locking 
A summary of the status of the auxiliary arm locking effort.

To help with lock acquisition we are attempting to independently lock the Y arm using light injected through ETMY. At present this secondary light source is an NPRO laser situated on the SP table. The laser light is transported to the ETM using a single mode optical fibre. In the future we might pick off some PSL light and apply a frequency shift.

We have been able to successfully mode match the fibre beam into the cavity and have been attempting lock the cavity using standard PDH signals (phase modulation sidebands are added to the light before it enters the fibre).

As yet no acceptable error signals have been produced. The demodulated RF signal is showing a time varying, bipolar dc offset.

We have minimised the residual amplitude modulation of the EOM but we expect small signals due to the undercoupled nature of the system, it could be that whatever RFAM still present is varying with time and causing this behaviour. We are also able to produce similar offsets by stressing (i.e. bending, shaking) the fibre. Could it be that the fibre is somehow converting PM into AM? Are we seeing etalon effects in the fibre or elsewhere?

If we cannot make any further progress with the existing setup we shall move the NPRO to the ETM table and try again. We are also looking into purchasing some other types of fibre.

Other things to consider are injecting through POY or using some other wavelength - neither seems obviously better.

Fiber, behavior
ELOG V3.1.3-